# Neural network prediction of model parameters for strong lensing samples from Hyper Suprime-Cam Survey

### v1

@@ 4 Results and Discussion
We train our network using simulated lenses and their SIE lens model parameters Einstein radius, axis ratio and position angle (see Appendix A for conventions) and apply it to simulated as well as real SuGOHI lenses to estimate the same parameters. Amongst the three parameters Einstein radius is the most crucial parameter to describe the lens mass distribution and it is also the most robust parameter to predict as it mostly depends on just the radial distances of the lensed arcs from the central lens galaxy. However, axis ratio and position angle are related to the actual configuration of the lensed arcs and are tricky to infer. We test the network on the simulated lenses as we have true model parameters for comparing the network predictions. In addition as mentioned previously, we also test on few of the SuGOHI lenses that have been already modelled in the literature. It was a challenge to find reliable parameters for these real lenses to compare the predictions from our network. To address this issue, we run a traditional MCMC modelling pipeline YL on the SuGOHI lenses to obtain the lens model parameters and compare our network predictions with the estimations from YL.
@@
When we train any CNN on a training sample which contains images having specific features, the network learns those features during the training process and we expect it to perform well when tested on the images having similar features. As a result, the performance of our network on the real SuGOHI lenses hugely depends on how well our simulated training images resemble the real SuGOHI lenses. As discussed in Section 2.1, we tried different training samples with various degrees of realism and we briefly describe the results here. We split each sample of simulated lenses in a training and a test sample. After training our network on a particular simulated training sample, we test it on the simulated as well as the real SuGOHI test sample (lens light removed by YL).
The training sample a, where we convolve our simulated arcs with the PSF and add Poisson noise to them without adding any background noise is a very ideal case. When we train our network with this sample, it produces really good results on the simulated test sample but fails miserably on the real SuHOHI lenses due to missing features like background noise in the training data. In order to address this issue, we create the training sample b where we add Gaussian background noise to the training sample a. Training the network with the sample b improves the performance of the network on the real SuGOHI test images, however the results were still not satisfactory.
Therefore, we decide to produce sample c by adding the simulated arcs to the image cutouts of the respective HSC galaxy (central lens galaxy used to model the arcs). Although this sample resembles well with the real SuGOHI lenses (not processed by YL) having the real background noise along with the light from the central lens galaxy, when we train our network on this sample, it performs poorly on the real data when compared with the predictions from YL (see Appendix B). The reason for this failure is network’s inability to model the light coming from the central lens galaxy, which either outshines or contaminates the background lensed arcs in most of the cases. This necessitates modelling and subtracting the lens light before feeding the images to the network and we use YL pipeline for this purpose. The training sample d is produced by processing the sample c with YL, identically as we process the real SuGOHI lenses to subtract the lens light. When we train our network on the training sample d, surprisingly it fails on both the simulated as well as real lenses, especially in predicting the axis ratio and the position angle. We investigate this issue further by performing different sanity checks with YL. Our experiments show that processing the lenses via YL to remove light from the central lens often introduces uncertainties in the lensed configuration, particularly, for lenses with small Einstein radii, either by leaving a residue or over-subtracting the flux. Even though the Einstein radius is a robust parameter to infer, it seems that the axis ratio and the position angle are quite sensitive to these uncertainties.
Thus, we finally create a training sample with a technique that does not involve adding or subtracting the lens light but has real foregrounds and backgrounds. We add our simulated lensed arcs to selected cutouts from HSC with central empty regions as discussed while concluding the Section 2.1 and use this training sample to produce all of the results shown in this paper. We have also compared our results for a subset of the SuGOHI lenses with the analysis done in S19 and S23 as we describe in the following sections.
4.1 Performance on simulated and real lenses
The performance of our network on the simulated test sample is shown in Fig. 2. In this plot, we compare the lens model parameters estimated by our network with the true lens model parameters that we used to simulate these lenses. We can see from the sample image in Fig. 2 that these simulated lenses have noisy features similar to the real SuGOHI lenses as we have added the simulated arcs to real HSC cutouts as described in the subsection  2.1. Our network performs reasonably in predicting the lens model parameters for the simulated lenses. Our network is able to recover the Einstein radius really well, while the scatter in the axis ratio and position angle is also within acceptable limits considering the difficulty in estimating these two parameters. We have also tested our network on the real SuGOHI grade A+B lenses as shown in the Fig. 3 and on a subsample consisting of only grade A lenses as shown in the Fig. 4. In Fig. 3 and Fig. 4, we compare the lens model parameters estimated by our network with the parameters predicted by YL along with their uncertainties. We can see that the quality of results on both the A+B and grade A lenses is similar, suggesting no peculiar advantage in selecting just the grade A lenses. The predictions from our network and YL mostly agree within the acceptable limits for the Einstein radius, while there is a quite disagreement between the two for axis ratio and position angle.
The two main challenges we faced while testing our network on the real lenses is lens-light subtraction and obtaining the true parameters of real lenses to compare our network predictions with. Generally in a lens system, the central lens galaxy outshines the lensed galaxies in the background and blending can further make it difficult to identify the lensed arcs. So, this contamination caused by the central lens light prompts us to remove it before feeding the image to the network. While we have conveniently avoided adding the lens light in the simulated images, we don’t have this liberty for the real lenses and we had to find a method to remove it. Nevertheless, quite a few of these SuGOHI lenses have been rigorously modelled in the literature, making it difficult to obtain reliable benchmarks for a larger sample to compare our results with. To address both of these aforementioned issues, we decided to use YL pipeline with some minor changes suited for our analysis.
@@ 4.1 Performance on simulated and real lenses
A sample of SuGOHI lenses processed by YL is shown in the Fig. 5. YL first models the light from the foreground lens galaxy in the centre and removes it to improve the identification of the lensed arcs in the background as shown in the second column (lens-subtracted) of Fig. 5. It then distinguishes potential lensed arcs from the foreground objects as shown in the segmentation map in the third column (object detection) of Fig. 5. YL then fits the SIE model to the lensed arcs in the lens light subtracted image as shown in the fourth column (lens model) of Fig. 5. YL also estimates the lens model parameters which we have used as benchmarks to compare our results with as shown in the Fig. 3 and Fig. 4. The errors on these parameters are from the 68 percent credible intervals from the posterior distribution of the parameters given the lens subtracted image, and obtained using a Monte Carlo Markov Chain.
@@
We find that the predictions for Einstein radius by YL and our network are in good quantitative agreement, although the axis ratio and the position angle fair poorly. However, axis ratio and position angle are quite sensitive to the actual arc configuration. When we process the SuGOHI lenses using YL to remove the central lens light, sometimes the pipeline does not perform well either leaving a residue from the central lens light or modelling a part of the lensed source as a lens light and then subtracting it. As a result the lens light subtraction process leaves its imprints which contaminates the actual configuration of lensed arcs, further leading to erroneous predictions of the lens model parameters by YL. The huge uncertainties on the parameter estimation from YL shown by the green error bars is also a concern. Given these issues, one may question the reliability of the parameters obtained by YL. To address this we also judge the performance of our network on the real SuGOHI lenses obtained from other competing methods as described in the next subsection.
4.2 Comparison with S19 and S23
@@
For this purpose, we select 10 grade A SuGOHI lenses (see Fig. 5 and Table 1) that were also modelled by S19 and S23. In S19, they modelled a sample of 23 strong lenses from the constant mas (CMASS) sample of Baryon Oscillation Spectroscopic Survey (BOSS) galaxies for which HSC imaging data in g, r, i, z and y bands is available. They model the lens mass with a SIE profile running a MCMC code, using the software EMCEE (Foreman-Mackey et al., 2013).
@@
In S23, they applied a CNN to 31 grade A real galaxy-scale lenses from SuGOHI and compared their results with traditional, MCMC sampling-based models obtained from their pipelines GLEE & GLAD. They also compared the results obtained from GLEE and GLAD with the results presented in S19 for some lenses common in both the analysis as shown in the fig.3 of S23. While both of these methods agree well on the predictions of Einstein radius, they often differ significantly for the axis ratio and position angle (often beyond the quoted uncertainties). The fact that S19 uses SIE-only model, where GLEE and GLAD uses SIE+external shear model, alone is not enough to justify the discrepancy (see discussion in S23).
@@
We process this sample of 10 grade A SuGOHI lenses using YL as shown in Fig. 5 in order to remove the central lens light and foreground objects in order to feed it to our network. We compare predictions from our network with the traditional MCMC modelling results presented in S19 and S23 in Fig. 6. We convert the Einstein radius and position angle from the other two methods to our convention before comparing (see Appendix A). The errors in Fig. 6 are obtained by quantifying the scatter in the network predictions on a simulated test sample shown in Fig. 2.
Here, we first describe our inferences from the visual comparison of the lens models shown in S19 (see fig. 1 in Sonnenfeld et al., 2019), S23 (see fig. B.1 - fig. B.31 in Schuldt et al., 2023) and YL (see Fig. 5) [^1] Note: Since our network does not make predictions for all of the lens and source parameters yet, we cannot produce the equivalent “best-fit” model images for visual comparison.. Next, we give a quantitative comparison of the model parameters from this work (Fig. 6) and the other studies. The errorbars mentioned in the following text and as shown in the (Fig. 6) are 1 $\sigma$ errorbars from our network predictions and are obtained by quantifying the scatter in the network predictions on a simulated test sample shown in Fig. 2.
@@
    HSCJ015618-010747 : In this case, the models of YL, S19 and S23 (see their fig. B.1) are visually similar. The axis ratio, $q$, predicted by S19 is close to unity and poses difficulty in constraining the PA which deviates from other results. The $q$ values predicted by S23 and network are within the quoted errorbars. The predictions for PA are within the errorbars for YL, S23 and network, while the predictions for $\theta_{\rm E}$ are close enough for all the methods.
@@
    HSCJ020241-064611 : For this lens, the models of YL, S19 and S23 (see their fig. B.3) look qualitatively similar with one image in the north and the other in the south direction. The model parameters $\theta_{\rm E}$ and $q$, for all four methods, are roughly consistent within the quoted errorbars due to the similarity of models containing two compact images, whereas the PAs of the mass distribution are harder to constrain due to their low ellipticities (i.e. high $q$).
@@
    HSCJ021737-051329 : The best-fit models of this lens by YL, S19 and S23 (see their fig. B.5) are similar and agree well on $\theta_{\rm E}$. Our $\theta_{\rm E}$ shows deviation from other methods. The PA is robustly constrained as all the methods give strongly consistent predictions. The predicted axis ratios are comparable for S19 and YL but the other two methods are not consistent with this, making $q$ the least well-constrained parameter despite the apparent similarity in their mass models.
@@
    HSCJ022346-053418 : Since the counter image of the arc, if any, is barely visible in this lens, we expect that the degeneracies in various models will become more apparent here. The model images of YL, S19 and S23 (see their fig. B.6) are visually similar. The model parameters $\theta_{\rm E}$ and $q$, for all four methods, are roughly consistent within the quoted errorbars. The PAs from S19, S23 and network are also approximately consistent within the errorbars, while the PA from YL deviates from this.
@@
    HSCJ022610-042011: In this case, the lens produces two images of the source galaxy. The model parameter $\theta_{\rm E}$, for all four methods, is consistent. There are differences in the predicted values of the axis ratio of the lensing potential, although all predictions are $>0.5$. We find that the axis ratio predicted by YL and S23 (for model, see their fig. B.7) agree well while those from S19 and network and are roughly within the errorbars. The PAs from S23 and network are within the quoted errorbars, while YL and S19 predict PA values which are close. The overall circularity ($q>0.5$) of the system implies that the position angle is hard to constrain and that can be seen from the large mismatch of the PAs inferred from the different methods as shown in Fig. 6.
@@
    HSCJ085855-010208 : The lens subtracted image in this lens system shows a near Einstein ring which is typical for a system with axial symmetry and a perfect alignment of the source with the lens center. The models from YL, S19 and S23 (see their fig. B.11) all reflect this feature resulting in values of the axis ratios close to unity for the lens model. The radius of the near perfect Einstein ring further helps each of the models to constrain the parameter $\theta_{\rm E}$ in a consistent manner. However, the inference of the PA is quite uncertain given the axial symmetry.
@@
    HSCJ121052-011905 : In this case, we can see the 3 images almost making an Einstein ring but two of them (the north-east and the west images) have extended structures deviating from the tangential direction. The models of YL and S19 are similar although the source of S19 is more compact than that of YL. These models lead to more smooth and circular configuration than the actual system. The S23 (see their fig. B.15) source model is more clumpy and does not have extended features. The inferred $\theta_{\rm E}$ agree with each other because of the similar angular separations of the arcs from the center of the lens potential. The model parameter $q$, predicted by all four methods is consistent within the errorbars. The PAs from YL, S23 and network are within the quoted errorbars, while the PA from S19 deviates from this.
@@
    HSCJ142720+001916 : For this lens, S19 and S23 (see their fig. B.22) models appear more accurate with an extended source whereas YL model is more circular with a compact source. The models from all methods agree well on $\theta_{\rm E}$ and the predictions for $q$ are also within the errorbar from our network. The inferred PA from S19 and S23 are distinct from both YL and our network, where the latter two are more consistent with each other.
@@
    HSCJ223733+005015 : In this case, YL and S19 models contain two diametrically opposite images and these models are qualitatively similar and accurate. As a result, these two methods also agree on the predictions of all the three parameters. The model of S23 (see their fig. B.27) does not seem correct with an arc modelled incorrectly in the south-east direction. However, the results of S23 for $\theta_{\rm E}$ and PA are within the errorbars from the network.
@@
    HSCJ230335+003703 : For this quad lens system, S19 model looks quite accurate with distinct four images of a compact source. The lens model of YL does not seem accurate. It has an extended source and the model looks more circular than the actual configuration of the lensed images. The model of S23 (see their fig. B.28) also does not do justice to the actual configuration.The inferred PAs from YL, S19 and S23 agree well and are within the errrorbars from the network. The $\theta_{\rm E}$ for all the methods is roughly consistent within the quoted errorbars. The axis ratio is not constrained well for this lens.

### v2

@@ 4 Results and Discussion
We train our network using simulated lenses and their true SIE lens model parameters, namely, the Einstein radius, axis ratio and position angle (see Appendix A for conventions). We further convert the axis ratio and the position angle into the two components of the ellipticity ($e_{x}$ and $e_{y}$) as discussed in the Appendix A. Amongst the three parameters Einstein radius is the most robustly predicted parameter as it depends on the radial angular separations of the lensed arcs from the central lens galaxy. However, the axis ratio and position angle are related to lensed arc properties such as their configuration, their orientation and their multiplicity and thus, are difficult to infer accurately. Presence of the lens light may further worsen their prediction.
We train our network with four different simulated samples with increasing realism. The PureSims sample, where we convolve our simulated arcs with the PSF and add Poisson noise to them without adding any background noise, is a very ideal case. In this case we expect the network to be thrown off due to the presence of background galaxies, noise fluctuations in real images or imperfect subtraction of the lens light. In order to address these issues incrementally, we have used additional training samples. In the GauNoise training sample, we generate Gaussian background noise equivalent to our expectation in the Subaru HSC survey and add it to the PureSims sample, such that the network learns about the presence of noise and becomes insensitive to noise peaks. The HSCempty training sample, comprises of lensed source images added to cutouts from real HSC data which do not contain bright objects near the centre, and thus contains both noise and unassociated galaxies in the cutout. Finally, we also create a LensLight training sample, where the simulated lensed sources are added to the HSC image of the galaxy which was used to model the gravitational lensed images, in order to test whether lens light can further provide useful information for parameter estimation.
We test the performance of the network trained on each of these sample on the corresponding simulated test lens sample as we have true model parameters for comparing the network predictions and we also have statistically large lens samples to work with. The results of this exercise are shown in the upper panels of Figs. 25. We quantify the predictions corresponding to these simulated test samples in a statistical manner. We divide the range of true values of each parameter in 10 bins, logarithmically for $\theta_{\rm E}$ and linearly for $e_{x}$ and $e_{y}$.
[…]
The performance of our network on PureSims, GauNoise, HSCempty and LensLight test samples can be compared using the statistical quantities defined above and the standard deviation ($\sigma$), bias ($b$) and outlier fraction ($f_{\rm{out}}$) are shown in the three different rows of Fig. 6, respectively. On the other hand, the three columns correspond to the parameters $\theta_{\rm E}$, $e_{x}$ and $e_{y}$, respectively.
For the Einstein radius, the PureSims trained network results in the lowest standard deviation in its predictions. However, this comes at the cost of an increasing bias as a function of increasing Einstein radius. The HSCempty sample has the lowest bias for the Einstein radius, although a larger scatter in the predictions. The outlier fraction estimates when trained on these network samples are quite noisy but seem to lie above the 5 percent thresholds in most cases, signifying the presence of non-Gaussian scatter in the predictions. The outlier fraction seems to be the lowest for the LensLight sample, again at the expense of an increasing bias at large Einstein radius. Overall the network trained on the GauNoise sample seems to represent a compromise in terms of the bias and scatter in the Einstein radius.
For the two ellipticity components, shown in the second and third column of the Fig. 6, respectively, the network trained on the LensLight sample results in the best results with the lowest scatter and bias, while the outlier fraction for all the samples are similar and hover around values between $5-12\%$. However, it is important to remember that the LensLight sample is simulated by keeping the components of ellipticity for the lens mass to be the same as the components of ellipticity for the lens light. Thus it is possible that the network is learning the components of ellipticity for the lens mass from the higher signal-to-noise ratio light from the lens present in the images, rather than from the lensed source configuration, itself. This also supports the presence of a larger bias and scatter in the Einstein radius prediction.
In order to further find support for this scenario, we simulated samples similar to the LensLight sample but introducing an offset between the ellipticity components of the lens potential and the light distribution. First, we add a randomly drawn Gaussian offset with a scatter of $\sigma_{\rm q}=0.05$ and $\sigma_{\rm PA}=10$°for the axis ratio and the position angle of the lens light, respectively, to obtain the lens potential axis ratio and position angle. These offsets are identical to those used by S21. We further split each of these samples in training and test samples and analyze with our network. The predicted ellipticity components even in the case of these offsets are very tightly correlated with the lens potential ellipticity due to the small offset.
We further increased the width of the Gaussian offsets in the axis ratio and PA to $0.1$ and $30$°, to test the extent to which these conclusions are valid. In Fig. 7, we show the predicted ellipticity components with respect to the lens light and the lens potential in the top and bottom row panels for the second sample, respectively, for a network trained on the LensLight sample with increased offsets. We see that the network gives ellipticity components which are closer to the values of the light from the lens, than the lens potential. This shows that if the lens potential and the lens light are known to have fairly small offsets, then the network trained on LensLight can be quite useful to infer this ellipticity. However, in the presence of large offsets, the network may not retrieve the lens potential ellipticity components.
@@
In addition to the simulated test samples, we also test our network trained on the different scenarios on the real SuGOHI lenses to assess the robustness of our network. When we train any CNN on a training sample which contains images having specific features, the network learns those features during the training process and we expect it to perform well when tested on the images having similar features. As a result, the performance of our network on the real SuGOHI lenses hugely depends on how well our simulated training images resemble the real SuGOHI lenses.
The two main challenges we faced while carrying out this comparison is lens-light subtraction as well as the estimates of the true parameters of real lenses to compare our network predictions with. For this purpose, we use the YL pipeline with some minor changes suited for our analysis, to remove the lens light, identify lensed features, and run a traditional MCMC modelling pipeline YL on the SuGOHI lenses to obtain the lens model parameters and compare them with our network predictions.
@@
A sample of SuGOHI lenses processed by YL is shown in the Fig. 8. YL first models the light from the foreground lens galaxy in the centre and removes it to improve the identification of the lensed arcs in the background as shown in the second column (lens-subtracted) of Fig. 8. It then distinguishes potential lensed arcs from the foreground objects as shown in the segmentation map in the third column (object detection) of Fig. 8. YL then fits the SIE model to the lensed arcs in the lens light subtracted image as shown in the fourth column (lens model) of Fig. 8. YL also estimates the lens model parameters which we have used to test our network.
The performance on the real SuGOHI lenses of our network when trained on the different samples is shown in the bottom panels of Figs. 2 -  5. Even though the network trained on PureSims gave the least scatter on the simulated test sample, it fares very poorly on the real SuGOHI lenses, likely due to the missing features in the training such as background noise and the presence of other galaxies. In particular, the network prediction for the Einstein radius shows no correlation with the Einstein radius of the real lenses. We believe that this is likely a result of the network latching on to either noise or other unassociated galaxies in the image as being a part of the lensed images of the source or due to imperfect subtraction of the lens light from the real SuGOHI lenses.
The network trained on the additional training samples help address these issues one by one. The performance on the real SuGOHI lenses of the network trained on the GauNoise sample is shown in the bottom panels of Fig. 3. Using a training sample which includes Gaussian noise improves the performance of the network on the real SuGOHI test images, where a positive correlation emerges in the network predictions of the Einstein radius compared to those from YL .
The performance on the real SuGOHI lenses of the network trained on the Lenslight sample is shown in the bottom panel of Fig. 5. When compared to the results from the GauNoise sample, we see lesser correlation in the prediction of the Einstein radius on the real SuGOHI sample, and an almost flat prediction of the ellipticity components. Although this sample resembles well with the realistic scenario, as it has the lens light, the noise as well as other unassociated galaxies, it appears that the network is unable to adequately separate the light coming from the central lens galaxy, which either outshines or contaminates the background lensed arcs due to blending. This shows the importance of accurate modelling and subtraction of the lens light before the images are fed to the network.
The results for the network trained using the HSCempty sample, which has all the desirable features like HSC foregrounds and backgrounds but no lens light to contaminate the arcs, on the real SuGOHI lenses are shown in the bottom panel of Fig. 4. We present the corresponding lens model parameters estimated by the network and YL , along with their errors, in the Table 2 (see Appendix D). Training on such samples improves the results for the SuGOHI test sample than when trained on the Lenslight sample and a performance comparable to the GauNoise sample.
Finally, we have also carried out cross-tests of the network trained on GauNoise sample with test lenses from HSCempty, and vice versa. Given that real lens galaxies are expected to have features such as background galaxies and real noise, we believe this is yet another important test, since the identification of lensed features by an alogorithm like YL may be susceptible to some failure modes. We find that the network trained on GauNoise performs significantly worse on all input parameters when tested on HSCempty sample, while the network trained on HSCempty sample performs robustly even on the GauNoise test sample.
Given all these consideration, we choose the network trained on the HSCempty sample to be our best case scenario. Although the GauNoise sample performs at par with the HSCempty sample on the simulated as well as SuGOHI lenses, it is important to consider the robustness of the network trained on the HSCempty sample. The fact that the HSCempty sample perform comparably with GauNoise sample indicates that the network trained on the HSCempty sample has also learned the real noise and the foreground objects and can distinguish them from the lensed arcs while predicting the lens mass parameters, which is important during application to real lenses.
@@
We acknowledge that when we process the SuGOHI lenses using YL to remove the central lens light, sometimes the pipeline does not perform well either leaving a residue from the central lens light or modelling a part of the lensed source as a lens light and then subtracting it. As a result the lens light subtraction process can leave residuals which contaminate the actual configuration of lensed arcs, which can affect the predictions of the lens model parameters by YL . Given these issues, one may question the reliability of the parameters obtained by YL in order to test the network. Therefore, we also judge the performance of our network on a subset of real SuGOHI lenses which have been analysed previously by S19 and S23b and present this comparison in the next subsection.
4.1 Network performance comparison of 10 real SuGOHI lenses commonly modelled by S19 and S23b
@@
A sample of 23 strong lenses from the constant mass (CMASS) sample of Baryon Oscillation Spectroscopic Survey (BOSS) galaxies was modeled in S19 for which HSC imaging data in g, r, i, z and y bands was available. They model the lens mass with a SIE profile running a MCMC code, using the software EMCEE (Foreman-Mackey et al., 2013). S23b applied a CNN to 31 grade A real galaxy-scale lenses from SuGOHI and compared their results with traditional, MCMC sampling-based models obtained from their pipelines GLEE & GLAD. They also compared the results obtained from GLEE and GLAD with the results presented in S19 for some lenses common in both the analysis as shown in the fig.3 of S23b. We select 10 grade A SuGOHI lenses (see Fig. 8 and Table 1) that were commonly modelled by S19 and S23b.
@@
While both of these methods agree well on the predictions of Einstein radius, they often differ significantly for the components of ellipticity (often beyond the quoted uncertainties). The fact that S19 uses SIE-only model, where GLEE and GLAD uses SIE+external shear model, alone is not enough to justify the discrepancy (see discussion in S23b).
@@
We process this sample of 10 grade A SuGOHI lenses using YL as shown in Fig. 8 in order to remove the central lens light and foreground objects in order to feed it to our network that was trained on the HSCempty sample. We compare predictions from our network with the traditional MCMC modelling results presented in S19 and S23b in Fig. 9. We convert the Einstein radius and position angle from the other two methods to our convention. We further convert the axis ratio and the position angle into components of ellipticity before comparing (see Appendix A).
[…]
@@
    HSCJ015618-010747 : In this case, the models of YL , S19 and S23b (see their fig. B.1) are visually similar. The $e_{x}$ and $e_{y}$ values predicted by S23b, YL and network are within the quoted errorbars, while the predictions for $\theta_{\rm E}$ are close enough for all the methods. The ellipticity components predicted by S19 are close to zero and deviate from other results.
@@
    HSCJ020241-064611 : For this lens, the models of YL , S19 and S23b (see their fig. B.3) look qualitatively similar with one image in the north and the other in the south direction. The values of $\theta_{\rm E}$ for all four methods, are consistent within the quoted errorbars due to the similarity of models containing two compact images, whereas the ellipticity components are not constrained well enough .
@@
    HSCJ021737-051329 : The best-fit models of this lens by YL , S19 and S23b (see their fig. B.5) are similar and agree well on $\theta_{\rm E}$ and $e_{y}$ Our $\theta_{\rm E}$ and $e_{x}$ show deviation from other methods, while our $e_{y}$ agrees with them.
@@
    HSCJ022346-053418 : Since the counter image of the arc, if any, is barely visible in this lens, we expect that the degeneracies in various models will become more apparent here. The model images of YL , S19 and S23b (see their fig. B.6) are visually similar. The model parameters $\theta_{\rm E}$ and $e_{y}$, for all four methods, are roughly consistent within the quoted errorbars, whereas the $e_{x}$ values are poorly constrained.
@@
    HSCJ022610-042011: In this case, the lens produces two images of the source galaxy. The model parameters $\theta_{\rm E}$ and $e_{y}$, for all four methods, are constrained well, while there is a huge inconsistency in the $e_{x}$ predictions.
@@
    HSCJ085855-010208 : The lens subtracted image in this lens system shows a near Einstein ring which is typical for a system with axial symmetry and a perfect alignment of the source with the lens center. The radius of the near perfect Einstein ring further helps each of the models to predict the parameter $\theta_{\rm E}$ accurately. The ellipticity components are also well constrained for S19, YL and our network, while predictions from S23b seem to deviate a from this.
@@
    HSCJ121052-011905 : In this case, we can see the 3 images almost making an Einstein ring but two of them (the north-east and the west images) have extended structures deviating from the tangential direction. The models of YL and S19 are similar although the source of S19 is more compact than that of YL . These models lead to more smooth and circular configuration than the actual system. The S23b (see their fig. B.15) source model is more clumpy and does not have extended features. The inferred $\theta_{\rm E}$ values for all the methods are within the errorbars because of the similar angular separations of the arcs from the center of the lens potential, whereas the ellipticity components are not well constrained due to the differences in the overall configuration.
@@
    HSCJ142720+001916 : For this lens, S19 and S23b (see their fig. B.22) models appear more accurate with an extended source whereas YL model is more circular with a compact source. The models from all methods agree well on $\theta_{\rm E}$ and the predictions for $e_{x}$ are also within the errorbar from our network, while the parameter $e_{y}$ is quite poorly constrained.
@@
    HSCJ223733+005015 : In this case, YL and S19 models contain two diametrically opposite images and these models are qualitatively similar and accurate. As a result, these two methods also agree on the predictions of all the three parameters. The model of S23b (see their fig. B.27) does not seem correct with an arc modelled incorrectly in the south-east direction. However, the results of S23b for $\theta_{\rm E}$ and $e_{x}$ agree well with our network.
@@
    HSCJ230335+003703 : For this quad lens system, S19 model looks quite accurate with distinct four images of a compact source. The lens model of YL does not seem accurate. It has an extended source and the model looks more circular than the actual configuration of the lensed images. The model of S23b (see their fig. B.28) also does not do justice to the actual configuration. The $\theta_{\rm E}$ for all the methods is roughly consistent within the quoted errorbars. The inferred $e_{y}$ from YL , S19 and S23b agree well but our network deviates from this, whereas for $e_{x}$ the predictions from YL and S19 are within the errorbar from our network.